# SC-LoRA: Balancing Efficient Fine-tuning and Knowledge Preservation via Subspace-Constrained LoRA

### v2

@@ Abstract
Minrui Luo∗1,2, Fuhang Kuang∗2, Yu Wang3, Zirui Liu2, Tianxing He†2,1thanks: Equal contribution.thanks: Corresponding author. 1Shanghai Qi Zhi Institute 2Institute for Interdisciplinary Information Sciences, Tsinghua University 3Institute of Information Engineering, Chinese Academy of Sciences {luomr22,kfh22,liu-zr22}@mails.tsinghua.edu.cn; wangyu2002@iie.ac.cnhetianxing@mail.tsinghua.edu.cn
@@ 1 Introduction
Recent works have found that carefully designed initialization on LoRA adapters can solve these problems. (Meng et al., 2024) initializes LoRA adapters by parts of Singular Value Decomposition (SVD) of original weight $W_{0}$, leading to faster convergence and improved performance by encapsulating the most significant information stored in $W_{0}$. Later works (Yang et al., 2024; Paischer et al., 2024) initialize LoRA weights based on semantic information stored in the activations of each layer on the target fine-tuning dataset. These data-driven approaches successfully enhance the fine-tuning speed and performance. Towards catastrophic forgetting problem in LoRA fine-tuning, (Yang et al., 2024) proposes to initialize LoRA weights by the least principal directions of world knowledge data features, successfully alleviating the forgetting problem. However, these works can only solve either side of the two problems, but do not consider the trade-off between enhancing fine-tuning performance and preserving pre-trained knowledge, which is a common need when doing parameter-efficient fine-tuning.
@@ 2 Related Work
PiSSA (Meng et al., 2024) argued that the default initialization of “Gaussian noise (He et al., 2015) and zero” to the adapters can lead to slow convergence. Hence they propose to apply singular value decomposition to original weight matrices and utilizes the top components to initialize LoRA, encapsulating the most significant information stored in original weights. CorDA (Yang et al., 2024) utilizes covariance matrices of data context, and takes the first (or last) singular vectors after context-oriented decomposition as initialization of LoRA adapters. They propose two different modes, one for improving fine-tuning performance and the other for mitigating world knowledge forgetting. Similar to CorDA, EVA (Paischer et al., 2024) feeds fine-tuning data into the model, applies sigular value decomposition to activation covariance matrices, and takes top singular vectors as initialization weights. LoRA-GA (Wang et al., 2024b) also utilizes data context but applies decomposition on the gradient. (Hayou et al., 2024) analyze the initialization of LoRA adapters, and have shown how the asymmetry of two low rank matrices affects training dynamics.
@@
Consequently, researchers have developed various defense strategies against such fine-tuning risks, generally falling into following approaches: enhancing the original safety alignment (Huang et al., 2024c, a), restricting the scope of fine-tuning parameters (Wei et al., 2024; Li et al., 2025), mixing additional safety data (Wang et al., 2024a; Huang et al., 2024b), modifying the loss function (Qi et al., 2025) and post-fine-tuning processing (Yia et al., 2024; Hsu et al., 2024). Different from previous works, our method focuses on an alternative approach of mitigating safety risks during fine-tuning by modifying initialization and model architecture, without mixing safety data during fine-tuning or appending prefix during inference time - both of which would incur computation overhead.
@@
Catastrophic forgetting (McCloskey and Cohen, 1989) is a phenomenon when models lose previously acquired knowledge when adapting to new tasks, and has been extensively studied in deep learning. Early approaches to solve the problem include knowledge distillation (Li and Hoiem, 2018), rehearsal (Riemer et al., 2019) and dynamic architectures (Yan et al., 2021). For large language models, preserving world knowledge remains challenging due to massive pre-training data and model size. Recent efforts mitigate forgetting by freezing pre-trained layers while introducing new adapters (Wu et al., 2024; Dou et al., 2024). Recently (Yang et al., 2024) proposed CorDA with Knowledge-Preserved Adaptation (KPA) mode, addressing world knowledge forgetting through LoRA initialization.
@@ 3 Method
Figure 1: Comparison of LoRA with default Kaiming initialization and our proposed SC-LoRA. (a) LoRA initializes down-projection matrix A by Gaussian noise and up-projection matrix B by zero matrix. (b) Our SC-LoRA initializes A by $Q_{r}^{\top}W_{0}$ and B by $Q_{r}$, where $Q_{r}$ consists of $r$ orthonormal vectors as columns obtained by Algorithm 1.
@@ 3.2 SC-LoRA
Known as catastrophic forgetting problem Chen et al. (2020), a large language model often performs worse on its pre-trained knowledge after fine-tuning on a downstream task. To this end, we consider fine-tuning a large language model on downstream task $T_{+}$, while preserving its ability on the other task $T_{-}$. Consider the output of a linear layer $h=W_{0}x=W_{\mathrm{res}}x+B_{\mathrm{init}}A_{\mathrm{init}}x$. We denote $\mathcal{P}_{+}$ and $\mathcal{P}_{-}$ the distribution of $h$ when the model is fed with data from $T_{+}$ and $T_{-}$, respectively. Our aim is to initialize $A,B$ within the $r$-rank constraint so that $BAx$ preserves the most of $\mathcal{P}_{+}$ and the least of $\mathcal{P}_{-}$, so that after initialization, the trainable term $BAx$ is constrained to primarily focus on $\mathcal{P}_{+}$ while avoiding modifying $\mathcal{P}_{-}$. This is equivalent to identify a low-dimensional subspace $S\subset\mathbb{R}^{d_{\mathrm{out}}}$ with rank $r$, on which the projection of $\mathcal{P}_{+}$ is mostly preserved and the projection of $\mathcal{P}_{-}$ is mostly eliminated. To evaluate such property of subspace $S$, we define the following reward:
@@
where $\beta\in[0,1]$ is a hyperparameter to tune. Here $\Pi_{S}:\mathbb{R}^{d_{\mathrm{out}}}\to S$ denote the orthogonal projection operator onto $S$. See Appendix A.1 for mathematical definition of $\Pi_{S}$.
@@
Let $h,x$ be the output and input of the original linear layer $W_{0}$, satisfying $h=W_{0}x$. When $A,B$ are initialized by Equations 7, 8, the following property holds:
@@
1:  Let $B_{+}=|\mathcal{D}_{+}|,B_{-}=|\mathcal{D}_{-}|$.
2:  Separately feed samples in $\mathcal{D}_{+},\mathcal{D}_{-}$ into the pre-trained model, collect batched output $\hat{X}_{+}\in\mathbb{R}^{d_{out}\times B_{+}},\hat{X}_{-}\in\mathbb{R}^{d_{out}\times B_{-}}$ of each linear layer. Within each sample, the output vector is summed over all tokens.
3:  $\mathrm{Cov}_{+}\leftarrow\frac{1}{B_{+}}\hat{X}_{+}\hat{X}_{+}^{\top}$.
4:  $\mathrm{Cov}_{-}\leftarrow\frac{1}{B_{-}}\hat{X}_{-}\hat{X}_{-}^{\top}$.
@@
6:  $Q_{r}\leftarrow(q_{1}\,q_{2}\,\cdots\,q_{r})$.
7:  $B_{\mathrm{init}}\leftarrow Q_{r}$.
8:  $A_{\mathrm{init}}\leftarrow Q_{r}^{\top}W_{0}$
9:  $W_{\mathrm{res}}\leftarrow W_{0}-B_{\mathrm{init}}A_{\mathrm{init}}$.
@@ 4 Experiments
For initialization of CorDA IPA and KPA mode, we calculate the covariance matrices with 256 samples from fine-tuning dataset and preserved knowledge dataset, respectively with 256 samples. We use AdamW optimizer Loshchilov and Hutter (2019) with the following hyper-parameters: batch size 128, learning rate 2e-5 (except for experiment in Section 4.2, where we tune the learning rate of baselines for better performance), cosine annealing learning rate schedule, warm-up ratio 0.03, and no weight decay. The rank of LoRA and its variants are all set to 128 for comparison. For SC-LoRA, we tune the hyperparameter $\beta$ to find a good balanced result. All experiment results are obtained by running on only one seed.
Below we discuss results in three settings: (1) Preservation of safety when fine-tuning on benign data; (2) Preservation of safety when fine-tuning on poisoned data; (3) Preservation of world knowledge when fine-tuning on math task.
@@ 4.1 Safety Preservation on Benign Finetuning
(Qi et al., 2024) has shown that fine-tuning on benign data can compromise the safety of aligned LLMs. In this setting, we aim to preserve the safety of aligned LLM while providing efficient fine-tuning on downstream tasks. Following the experimental settings by (Qi et al., 2025), we fine-tune Llama-2-7b-Chat model with safety alignment (Touvron et al., 2023) on Samsum (Gliwa et al., 2019) for 1 epoch. Samsum is a dataset for conversation summarization task, containing 14732 training samples.
To initialize our SC-LoRA model, we randomly select 256 samples from fine-tuning dataset ($\mathcal{D}_{+}$) Samsum to compute covariance matrix $\mathrm{Cov}_{+}$ for each linear layer, then use 256 harmful-question&refusal-answer pairs (as the safety dataset $\mathcal{D}_{-}$) provided by (Qi et al., 2025) to compute $\mathrm{Cov}_{-}$. These two collections of samples are also used to compute the covariance matrices of CorDA IPA and CorDA KPA respectively.
For utility evaluation, we employ the standard ROUGE-1 score (Lin, 2004) for Samsum. For safety evaluation, we let the fine-tuned models to generate answers for 330 malicious questions provided by (Qi et al., 2024) (distinct from malicious questions for initialization) and employ DeepSeek-V3 (DeepSeek-AI et al., 2025) API to judge the harmfulness, assigning each answer an integer score from 1 (safe) to 5 (most harmful). We report the average score as harmfulness score of the model and the fraction of maximum-risk responses (score = 5) as harmfulness rate. Lower values for both metrics indicate stronger safety of the model.
@@ 4.2 Safety Preservation on Data Poisoning Attack
Harmful data injection is a common attack method to degrade the safety of LLMs during fine-tuning (Huang et al., 2024a, b, c). In this experiment, we aim to preserve safety in poisoned data scenarios. To construct the poisoned dataset, we first take 25600 data samples of MetaMathQA (Yu et al., 2024), then replace 1% of the data by harmful question-answer pairs provided by (Qi et al., 2024). We train each method for 1 epoch on the poisoned dataset. We use 256 samples from MetaMathQA for the initialization of SC-LoRA and CorDA IPA. The safety samples used for the initialization of SC-LoRA and CorDA KPA are the same with the previous experiment (Section 4.1).
For utility evaluation, we compute the answer accuracy on the validation set of GSM8k (Cobbe et al., 2021). Safety evaluation follows the setting in the previous section 4.1. For better comparability, we tune the learning rate of LoRA to 2e-5, 5e-5 and 1e-4. The learning rate for other methods is fixed to 2e-5.
@@
[…]
Llama-2-7b - 52.52 18.86 5.86 25.75 - - -
Full fine-tuning 6738M 47.42 4.16 6.64 19.41 50.27 6.94 28.60
LoRA 320M 46.81 1.05 7.04 18.30 41.77 5.46 23.62
PiSSA 320M 47.44 3.32 6.84 19.20 51.63 7.70 29.67
CorDA IPA 320M 30.20 9.83 5.41 15.15 51.40 8.34 29.87
CorDA KPA 320M 46.21 10.64 7.33 21.39 45.03 6.54 25.79
SC-LoRA $\beta=0$ 320M 44.26 5.18 7.19 18.88 53.53 8.98 31.25
$\beta=0.5$ 320M 48.91 7.70 6.89 21.17 53.37 8.62 31.00
$\beta=0.8$ 320M 50.52 10.64 7.04 22.73 52.46 7.62 30.04
4.3 World Knowledge Preservation
Pre-trained LLMs also have other pre-trained knowledge that is easy to lose after fine-tuning on downstream tasks, such as world knowledge Yang et al. (2024). In this setting, we aim to preserve the intrinsic world knowledge (e.g., common sense) within the pre-trained LLM while providing efficient fine-tuning on downstream tasks. We fine-tune the Llama-2-7b model (Touvron et al., 2023) on math task and evaluate its math ability (utility) and world knowledge performance. We train on 100000 samples of MetaMathQA (Yu et al., 2024) for 1 epoch and evaluate its math ability on GSM8k (Cobbe et al., 2021) and MATH (Yu et al., 2024) validation sets. World knowledge is evaluated by the exact matching score on TriviaQA (Joshi et al., 2017), NQ-open (Lee et al., 2019), and WebQS (Berant et al., 2013) through Evaluation-Harness Gao et al. (2024). We select 256 random samples from NQ-open as world knowledge samples used for the initialization of SC-LoRA and CorDA KPA, and 256 random samples from MetaMathQA as fine-tuning dataset for initializing SC-LoRA and CorDA IPA.
@@ 4.3 World Knowledge Preservation
4.4 Experimental Analysis on the Functionality of Hyper-Parameter $\beta$

### v3

@@ Abstract
Minrui Luo∗1,2, Fuhang Kuang∗1, Yu Wang3, Zirui Liu1, Tianxing He†2,1,4thanks: Equal contribution.thanks: Tianxing He is the corresponding author. 1Shanghai Qi Zhi Institute 2Institute for Interdisciplinary Information Sciences, Tsinghua University 3Institute of Information Engineering, Chinese Academy of Sciences 4Xiongan AI Institute {luomr22,kfh22,liu-zr22}@mails.tsinghua.edu.cn; wangyu2002@iie.ac.cnhetianxing@mail.tsinghua.edu.cn
@@ 1 Introduction
Recent works have found that carefully designed initialization on LoRA adapters can solve these problems. Meng et al. (2024) initializes LoRA adapters by parts of Singular Value Decomposition (SVD) of original weight $W_{0}$, leading to faster convergence and improved performance by encapsulating the most significant information stored in $W_{0}$. Later works (Yang et al., 2024; Paischer et al., 2024) initialize LoRA weights based on semantic information stored in the activations of each layer on the target fine-tuning dataset. These data-driven approaches successfully enhance the fine-tuning speed and performance. Towards catastrophic forgetting problem in LoRA fine-tuning, Yang et al. (2024) proposes to initialize LoRA weights by the least principal directions of world knowledge data features, successfully alleviating the forgetting problem. However, these works can only solve either side of the two problems, but do not consider the trade-off between enhancing fine-tuning performance and preserving pre-trained knowledge, which is a common need when doing parameter-efficient fine-tuning.
@@ 2 Related Work
PiSSA (Meng et al., 2024) argued that the default initialization of “Gaussian noise (He et al., 2015) and zero” to the adapters can lead to slow convergence. Hence they propose to apply singular value decomposition to original weight matrices and utilizes the top components to initialize LoRA, encapsulating the most significant information stored in original weights. CorDA (Yang et al., 2024) utilizes covariance matrices of data context, and takes the first (or last) singular vectors after context-oriented decomposition as initialization of LoRA adapters. They propose two different modes, one for improving fine-tuning performance and the other for mitigating world knowledge forgetting. Similar to CorDA, EVA (Paischer et al., 2024) feeds fine-tuning data into the model, applies sigular value decomposition to activation covariance matrices, and takes top singular vectors as initialization weights. LoRA-GA (Wang et al., 2024b) also utilizes data context but applies decomposition on the gradient. Hayou et al. (2024) analyze the initialization of LoRA adapters, and have shown how the asymmetry of two low rank matrices affects training dynamics.
@@
Consequently, researchers have developed various defense strategies against such fine-tuning risks, generally falling into following approaches: enhancing the original safety alignment (Huang et al., 2024c, a; Li et al., 2025a), restricting the gradient of fine-tuning parameters or the scope of trained residuals (Wei et al., 2024; Li et al., 2025b), mixing additional safety data (Wang et al., 2024a; Huang et al., 2024b), modifying the loss function (Qi et al., 2025) and post-fine-tuning processing (Yia et al., 2024; Hsu et al., 2024). Different from previous works, our method focuses on an alternative approach of mitigating safety risks during fine-tuning by only modifying initialization, without mixing safety data during fine-tuning, appending prefix during inference time, or adding extra high-rank modules to the model - which would incur computation overhead either in training or inference time.
@@
Catastrophic forgetting (McCloskey and Cohen, 1989) is a phenomenon when models lose previously acquired knowledge when adapting to new tasks, and has been extensively studied in deep learning. Early approaches to solve the problem include knowledge distillation (Li and Hoiem, 2018), rehearsal (Riemer et al., 2019) and dynamic architectures (Yan et al., 2021). For large language models, preserving world knowledge remains challenging due to massive pre-training data and model size. Recent efforts mitigate forgetting by freezing pre-trained layers while introducing new adapters (Wu et al., 2024; Dou et al., 2024). Recently Yang et al. (2024) proposed CorDA with Knowledge-Preserved Adaptation (KPA) mode, addressing world knowledge forgetting through LoRA initialization.
@@ 3 Method
Figure 1: Comparison of LoRA with default Kaiming initialization and our proposed SC-LoRA. (a) LoRA initializes down-projection matrix $A$ by Gaussian noise and up-projection matrix $B$ by zero matrix. (b) Our SC-LoRA initializes AA by $Q_{r}^{\top}W_{0}$ and B by $Q_{r}$, where $Q_{r}$ consists of $r$ orthonormal vectors as columns obtained by Algorithm 1.
@@ 3.2 SC-LoRA
Known as catastrophic forgetting problem Chen et al. (2020), a large language model often performs worse on its pre-trained knowledge after fine-tuning on a downstream task. To this end, we consider fine-tuning a large language model on downstream task $T_{+}$, while preserving its ability on the other task $T_{-}$. Consider the output of a linear layer $h=W_{0}x=W_{\mathrm{res}}x+B_{\mathrm{init}}A_{\mathrm{init}}x$. We denote $\mathcal{P}_{+}$ and $\mathcal{P}_{-}$ the distribution of $h$ when the model is fed with data from $T_{+}$ and $T_{-}$, respectively. Our aim is to initialize $A,B$ within the $r$-rank constraint so that $BAx$ preserves the most of $\mathcal{P}_{+}$ and the least of $\mathcal{P}_{-}$, so that after initialization, the trainable term $BAx$ is constrained to primarily focus on $\mathcal{P}_{+}$ while avoiding modifying $\mathcal{P}_{-}$. This is equivalent to identify a low-dimensional subspace $S\subset\mathbb{R}^{d_{\mathrm{out}}}$ with rank $r$, on which the projection of $\mathcal{P}_{+}$ is mostly preserved and the projection of $\mathcal{P}_{-}$ is mostly eliminated. To evaluate such property of subspace $S$, we define the following projection $\Pi_{S}$ and reward $R(S)$:
Definition 1.
$\Pi_{S}(x)$ $=\sum_{i=1}^{r}(q_{i}^{\top}x)q_{i}$ (1)
$=\sum_{i=1}^{r}(q_{i}q_{i}^{\top})x,\forall x\in\mathbb{R}^{n}.$
Note: the selection of the orthonormal basis does not affect $\Pi_{S}$.
@@
where $\beta\in[0,1]$ is a hyperparameter to tune.
@@
Let $h,x$ be the output and input of the original linear layer $W_{0}$, satisfying $h=W_{0}x$. When $A,B$ are initialized by Equations 8, 9, the following property holds:
@@
1: Let $B_{+}=|\mathcal{D}_{+}|,B_{-}=|\mathcal{D}_{-}|$, $L$ be the length of each sample (clipped to same length).
2: Separately feed samples in $\mathcal{D}_{+},\mathcal{D}_{-}$ into the pre-trained model, collect batched output $\hat{X}_{+}\in\mathbb{R}^{d_{out}\times B_{+}L},\hat{X}_{-}\in\mathbb{R}^{d_{out}\times B_{-}L}$ of each linear layer. Within each sample, the output vector is summed over all tokens.
[…]
@@
[…]
@@ 4 Experiments
For initialization of CorDA IPA and KPA mode, we calculate the covariance matrices with 256 samples from fine-tuning dataset and preserved knowledge dataset, respectively with 256 samples. We use AdamW optimizer (Loshchilov and Hutter, 2019) with the following hyper-parameters: batch size 128, learning rate 2e-5 (except for experiment in Section 4.3, where we tune the learning rate of baselines for better performance), cosine annealing learning rate schedule, warm-up ratio 0.03, and no weight decay. The rank of LoRA and its variants are all set to 128 for comparison. For SC-LoRA, we tune the hyperparameter $\beta$ to find a good balanced result. All experiment results are obtained by running on only one seed.
Below we discuss results in three settings: (1) Preservation of world knowledge when fine-tuning on math task; (2) Preservation of safety when fine-tuning on benign data; (3) Preservation of safety when fine-tuning on poisoned data.
[…]
Llama-2-7b - 52.52 18.86 5.86 25.75 - - -
Full fine-tuning 6738M 47.42 4.16 6.64 19.41 50.27 6.94 28.60
LoRA 320M 46.81 1.05 7.04 18.30 41.77 5.46 23.62
PiSSA 320M 47.44 3.32 6.84 19.20 51.63 7.70 29.67
CorDA IPA 320M 30.20 9.83 5.41 15.15 51.40 8.34 29.87
CorDA KPA 320M 46.21 10.64 7.33 21.39 45.03 6.54 25.79
SC-LoRA
$\beta=0$ 320M 44.26 5.18 7.19 18.88 53.53 8.98 31.25
$\beta=0.5$ 320M 48.91 7.70 6.89 21.17 53.37 8.62 31.00
$\beta=0.8$ 320M 50.52 10.64 7.04 22.73 52.46 7.62 30.04
4.1 World Knowledge Preservation
Pre-trained LLMs also have other pre-trained knowledge that is easy to lose after fine-tuning on downstream tasks, such as world knowledge Yang et al. (2024). In this setting, we aim to preserve the intrinsic world knowledge (e.g., common sense) within the pre-trained LLM while providing efficient fine-tuning on downstream tasks. We fine-tune the Llama-2-7b model (Touvron et al., 2023) on math task and evaluate its math ability (utility) and world knowledge performance. We train on 100000 samples of MetaMathQA (Yu et al., 2024) for 1 epoch and evaluate its math ability on GSM8k (Cobbe et al., 2021) and MATH (Yu et al., 2024) validation sets. World knowledge is evaluated by the exact matching score on TriviaQA (Joshi et al., 2017), NQ-open (Lee et al., 2019), and WebQS (Berant et al., 2013) through Evaluation-Harness Gao et al. (2024). We select 256 random samples from NQ-open as world knowledge samples used for the initialization of SC-LoRA and CorDA KPA, and 256 random samples from MetaMathQA as fine-tuning dataset for initializing SC-LoRA and CorDA IPA. Note that samples used in initialization are separate from those in evaluation.
@@ 4.2 Safety Preservation on Benign Finetuning
Qi et al. (2024) has shown that fine-tuning on benign data can compromise the safety of aligned LLMs. In this setting, we aim to preserve the safety of aligned LLM while providing efficient fine-tuning on downstream tasks. Following the experimental settings by Qi et al. (2025), we fine-tune Llama-2-7b-Chat model with safety alignment (Touvron et al., 2023) on Samsum (Gliwa et al., 2019) for 1 epoch. Samsum is a dataset for conversation summarization task, containing 14732 training samples and 819 testing samples.
To initialize our SC-LoRA model, we randomly select 256 samples from training set of Samsum ($\mathcal{D}_{+}$) to compute covariance matrix $\mathrm{Cov}_{+}$ for each linear layer, then use 256 harmful-question&refusal-answer pairs (as the safety dataset $\mathcal{D}_{-}$) provided by Qi et al. (2025) to compute $\mathrm{Cov}_{-}$. These two collections of samples are also used to compute the covariance matrices of CorDA IPA and CorDA KPA respectively. For better comparability, we tune the learning rate of all LoRA-based baselines to 2e-5, 5e-5 and 1e-4, and pick the best result among all learning rates. The learning rate for SC-LoRA is fixed to 2e-5. Full results of learning rate tuning can be seen in Appendix C Figure 5.
For utility evaluation, we employ the standard ROUGE-1 score (Lin, 2004) for testing set of Samsum. For safety evaluation, we let the fine-tuned models to generate answers for 330 malicious questions provided by Qi et al. (2024) (distinct from malicious questions for initialization) and employ DeepSeek-V3 (DeepSeek-AI et al., 2025) API to judge the harmfulness, assigning each answer an integer score from 1 (safe) to 5 (most harmful). We report the average score as harmfulness score of the model and the fraction of maximum-risk responses (score = 5) as harmfulness rate. Lower values for both metrics indicate stronger safety of the model.
@@
[…]
Data poisoning 256 1.136 1.212 45.26
128 1.155 2.121 45.64
64 1.200 2.727 46.10
32 1.864 1.616 45.72
Benign finetuning 256 1.097 0.000 51.67
128 1.118 1.515 51.91
64 1.145 1.515 52.46
32 1.158 1.515 52.42
Table 4: Sensitivity to the Number of Initialization Samples.
@@ 4.3 Safety Preservation on Data Poisoning Attack
Harmful data injection is a common attack method to degrade the safety of LLMs during fine-tuning (Huang et al., 2024a, b, c). In this experiment, we aim to preserve safety in poisoned data scenarios. To construct the poisoned dataset, we first take 25600 data samples from training set of MetaMathQA (Yu et al., 2024), then replace 1% of the data by harmful question-answer pairs provided by (Qi et al., 2024). We train each method for 1 epoch on the poisoned dataset. For the initialization of SC-LoRA and CorDA IPA, we use 256 samples from training set of MetaMathQA. The safety samples used for the initialization of SC-LoRA and CorDA KPA are the same with the previous experiment (Section 4.2).
For utility evaluation, we compute the answer accuracy on the validation set of GSM8k (Cobbe et al., 2021). Safety evaluation follows the setting in the previous section 4.2. For better comparability, we tune the learning rate of all LoRA-based baselines to 2e-5, 5e-5 and 1e-4, and pick the best result among all learning rates. The learning rate for SC-LoRA is fixed to 2e-5. Full results of learning rate tuning can be seen in Appendix C Figure 6.
@@
4.4 Ablation: Experimental Analysis on the Functionality of Hyper-Parameter $\beta$